OpenStudy (gudetamas):

Write an equation of the line passing through the point (5, 1) that is perpendicular to the line 5x+3y=15. (y=mxtb format)

OpenStudy (anonymous):

Perpendicular to a line means the negative reciprocal of the slope. So first, let's find the slope of x+7y=8. To do that, we need to rearrange it to slope intercept form (y=mx+b).

OpenStudy (solomonzelman):

If you have a know that 5x+3y=15 ─> y=(-5/3)x+5, then it's slope has to be \(-1/\color{red}{(-5/3)}\)

OpenStudy (anonymous):

Negative reciprocal of −17 is 7. So the perpendicular line has a slope of 7. And, we can begin to create the equation for the perpendicular line: y=7x+b. Using the information in the original problem, we can find the value of the y-intercept, b. The line must pass through (-4, 42). Simply plug those values into y=7x+b and solve for b. −42=7(−4)+b −42=−28+b −42+28=b −14=b

OpenStudy (gudetamas):

so y=7x-14

OpenStudy (wykeise):

Negative reciprocal of −17 is 7. So the perpendicular line has a slope of 7. And, we can begin to create the equation for the perpendicular line: y=7x+b. Using the information in the original problem, we can find the value of the y-intercept, b. The line must pass through (-4, 42). Simply plug those values into y=7x+b and solve for b. −42=7(−4)+b −42=−28+b −42+28=b −14=b

OpenStudy (alexandervonhumboldt2):

stop spamming

OpenStudy (alexandervonhumboldt2):

and negative reciprocal of -17 is 1/17

OpenStudy (alexandervonhumboldt2):

@Compassionate mods see deleted comments have fun in being rude!

OpenStudy (gudetamas):

@SolomonZelman also do know 5x+2y=12. in slope intercept form?

OpenStudy (gudetamas):

i just wanna know how to make 5x+2y=12. in slope intercept form jc

OpenStudy (anonymous):

yeah im back and i have a hole lot of them yea im talking to u alexander and this site manger the police

OpenStudy (anonymous):

u willy wonka

OpenStudy (alexandervonhumboldt2):

Please stop spaming

OpenStudy (anonymous):

Refer to the attachment created with the Mathematica program.

OpenStudy (gudetamas):

@robtobey ty